Q8. The Farmer's Market has $725,000 in sales. The profit margin is 4.1 percent and the
firm has 8,500 shares of stock outstanding. The market price per share is $23.
What is the price-earnings ratio?
can you show work
Answer #1

Profit margin=Net income/Sales

Net income=725000*4.1%=$29725

Earnings per share=29725/8500

=$3.49705882

P/E ratio=stock price/EPS

=(23/3.49705882)

=6.58(Approx)